# PAC-Bayesian Reinforcement Learning Trains Generalizable Policies

### v1

@@ 6 Conclusion
We introduced PB-SAC, a PAC-Bayesian actor-critic algorithm that bridges the gap between learning theory and practical deep RL by making practical use of non-vacuous performance certificates as active components of the learning process. Our theoretical contribution (a PAC-Bayesian bound with explicit mixing-time dependence and improved discount factor scaling) enables meaningful generalization guarantees in sequential decision-making. Algorithmically, our key innovations include policy-posterior synchronization, adaptive sampling curriculum for critic adaptation, posterior-guided exploration, and prior moving average that maintains competitive performance while providing formal certificates.
Several limitations warrant consideration. Our approach requires accurate mixing time estimation (underestimation leads to overconfident bounds), faces computational overhead that may limit scalability to large networks, and uses KL divergence which, despite its analytical convenience, is unstable when distributions diverge significantly and does not respect the geometry of the parameter space (Viallard et al., 2023). Future work could explore efficient posterior approximations, meaningful representations (layer/parameter-level rather than full network posteriors), adaptive mixing time estimation, Wasserstein distance alternatives to KL (Amit et al., 2022; Haddouche & Guedj, 2023; Viallard et al., 2023), and extensions beyond actor-critic frameworks. Despite these constraints, our results establish PAC-Bayesian RL as viable for certified performance in modern reinforcement learning, particularly valuable in high-stakes applications requiring theoretical guarantees.
@@ B.1 Bounded-differences property for MDP trajectories
As defined in equation 5, our empirical and expected losses are:
@@ B.2 Quantifying the impact of perturbed transitions
Suppose we replace a single transition at position $h$ in trajectory $j$. The change in the discounted return of this trajectory is bounded by:
$|G(\xi^{(j)})-G(\bar{\xi}^{(j)})|$ $=|\gamma^{h-1}(R_{h}-\bar{R}_{h})+\text{effects on future rewards}|$
$\leq\gamma^{h-1}R_{\max}+\text{effects on future rewards}$
Crucially, this perturbation affects not only the immediate reward but potentially all subsequent transitions and rewards in that trajectory. The change in our empirical loss is therefore bounded by:
@@ B.4 Full accounting of perturbation propagation effects
When divided by $T$ (because $\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)$ averages over $T$ trajectories), we get exactly the same quantity as the sum of coefficients above. Therefore, the bounded-differences condition holds with equality, confirming that our derivation of $\|c\|^{2}$ fully accounts for all propagation effects without requiring additional constants.
@@ B.5.1 From McDiarmid to moment generating function
McDiarmid’s inequality for Markov chains (equation 14) provides a concentration inequality on the deviation between empirical and expected losses. From this, we can derive a bound on the moment generating function (MGF) as shown by Paulin (2018):
@@
$\mathbb{E}_{\mathfrak{D}\sim\mathcal{M}^{(T)}}\left[\exp\left(\kappa(\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)-\mathcal{L}(\theta))\right)\right]\leq\exp\left(\frac{\kappa^{2}\|c\|^{2}\tau_{\min}}{8}\right)$ (15)
@@
where $\tau_{\min}$ is the mixing time of the Markov chain induced by policy $\pi_{\theta}$.
B.5.2 PAC-Bayes change of measure
Now we can follow the standard PAC-Bayes derivation. Let $\Theta$ be our parameter space and let $\mu\in\mathcal{P}(\Theta)$ be a prior distribution over $\Theta$ chosen independently of the data. For any posterior distribution $\rho\in\mathcal{P}(\Theta)$ (which may depend on $\mathfrak{D}$), we apply the change of measure inequality that follows from Donsker–Varadhan variational formula (Donsker & Varadhan, 1975; 1983).
Let $f(\theta)=\kappa(\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)-\mathcal{L}(\theta))$. Applying Lemma A.2:
@@ B.5.2 PAC-Bayes change of measure
$\mathbb{E}_{\theta\sim\rho}[\kappa(\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)-\mathcal{L}(\theta))]$ $\leq\mathrm{KL}(\rho\|\mu)+\ln\mathbb{E}_{\theta\sim\mu}[\exp(\kappa(\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)-\mathcal{L}(\theta)))]$ (16)
B.5.3 Combining with the MGF bound
@@
$\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}[\kappa(\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)-\mathcal{L}(\theta))]$ $\leq\mathrm{KL}(\rho\|\mu)+\mathbb{E}_{\mathfrak{D}}\ln\mathbb{E}_{\theta\sim\mu}[\exp(\kappa(\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)-\mathcal{L}(\theta)))]$ (17)
By Jensen’s inequality, since $\ln$ is concave:
$\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}[\kappa(\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)-\mathcal{L}(\theta))]$ $\leq\mathrm{KL}(\rho\|\mu)+\ln\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\mu}[\exp(\kappa(\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)-\mathcal{L}(\theta)))]$ (18)
By Fubini’s theorem (exchanging the order of expectations) and Lemma B.1:
$\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}[\kappa(\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)-\mathcal{L}(\theta))]$ $\leq\mathrm{KL}(\rho\|\mu)+\ln\mathbb{E}_{\theta\sim\mu}\mathbb{E}_{\mathfrak{D}}[\exp(\kappa(\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)-\mathcal{L}(\theta)))]$ (19)
$\leq\mathrm{KL}(\rho\|\mu)+\ln\mathbb{E}_{\theta\sim\mu}\left[\exp\left(\frac{\kappa^{2}\|c\|^{2}\tau_{\min}}{8}\right)\right]$ (20)
$=\mathrm{KL}(\rho\|\mu)+\frac{\kappa^{2}\|c\|^{2}\tau_{\min}}{8}$ (21)
@@
$\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}[\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)-\mathcal{L}(\theta)]$ $\leq\frac{\mathrm{KL}(\rho\|\mu)}{\kappa}+\frac{\kappa\|c\|^{2}\tau_{\min}}{8}$ (22)
B.5.4 High-probability bound via Markov’s inequality
Now, we convert this expectation bound into a high-probability bound. By Markov’s inequality A.1, for any non-negative random variable $X$ and $\delta>0$:
With probability at least $1-\delta$:
$\mathbb{E}_{\theta\sim\rho}[\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)-\mathcal{L}(\theta)]$ $\leq\frac{\mathrm{KL}(\rho\|\mu)+\ln\frac{2}{\delta}}{\kappa}+\frac{\kappa\|c\|^{2}\tau_{\min}}{8}$ (23)
B.5.5 Optimizing the bound
To tighten the bound, we minimize the right-hand side with respect to $\kappa>0$. Taking the derivative and setting it to zero:
$\frac{\partial}{\partial\kappa}\left(\frac{\mathrm{KL}(\rho\|\mu)+\ln\frac{2}{\delta}}{\kappa}+\frac{\kappa\|c\|^{2}\tau_{\min}}{8}\right)$ $=0$ (24)
$-\frac{\mathrm{KL}(\rho\|\mu)+\ln\frac{2}{\delta}}{\kappa^{2}}+\frac{\|c\|^{2}\tau_{\min}}{8}$ $=0$ (25)
Solving for the optimal $\kappa^{*}$:
$\kappa^{*}$ $=\sqrt{\frac{8(\mathrm{KL}(\rho\|\mu)+\ln\frac{2}{\delta})}{\|c\|^{2}\tau_{\min}}}$ (26)
Substituting $\kappa^{*}$ back into our bound:
$\mathbb{E}_{\theta\sim\rho}[\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)-\mathcal{L}(\theta)]$ $\leq\frac{\mathrm{KL}(\rho\|\mu)+\ln\frac{2}{\delta}}{\kappa^{*}}+\frac{\kappa^{*}\|c\|^{2}\tau_{\min}}{8}$ (27)
$=\sqrt{\frac{\|c\|^{2}\tau_{\min}(\mathrm{KL}(\rho\|\mu)+\ln\frac{2}{\delta})}{8}}+\sqrt{\frac{\|c\|^{2}\tau_{\min}(\mathrm{KL}(\rho\|\mu)+\ln\frac{2}{\delta})}{8}}$ (28)
$=\sqrt{\frac{\|c\|^{2}\tau_{\min}(\mathrm{KL}(\rho\|\mu)+\ln\frac{2}{\delta})}{2}}$ (29)
B.5.6 Final bound
Finally, substituting the expression for $\|c\|^{2}$ from equation 7:
$\mathbb{E}_{\theta\sim\rho}[\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)-\mathcal{L}(\theta)]$ $\leq\sqrt{\frac{\frac{R_{\max}^{2}}{T}\cdot\frac{1-\gamma^{2H}}{1-\gamma^{2}}\cdot\tau_{\min}\cdot(\mathrm{KL}(\rho\|\mu)+\ln\frac{2}{\delta})}{2}}$ (30)
$=\sqrt{\frac{R_{\max}^{2}\tau_{\min}(1-\gamma^{2H})}{2T(1-\gamma^{2})}\Big(\mathrm{KL}(\rho\|\mu)+\ln\frac{2}{\delta}\Big)}$ (31)
$\mathbb{E}_{\theta\sim\rho}[V_{\pi_{\theta}}]\geq\mathbb{E}_{\theta\sim\rho}[-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)]-\sqrt{\frac{R_{\max}^{2}\tau_{\min}(1-\gamma^{2H})}{2T(1-\gamma^{2})}\Big(\mathrm{KL}(\rho\|\mu)+\ln\frac{2}{\delta}\Big)}$ (32)
@@ B.5.6 Final bound
B.6 A note on the Markov assumption for Bellman errors

### v2

@@ 6 Conclusion
We introduced PB-SAC, a PAC-Bayesian soft actor-critic algorithm that bridges the gap between learning theory and practical deep RL by making practical use of non-vacuous performance certificates as active components of the learning process. Our theoretical contribution (a PAC-Bayes bound with explicit mixing-time dependence and improved discount factor scaling) enables meaningful generalization guarantees in sequential decision-making. Algorithmically, our key innovations include policy-posterior synchronization, adaptive sampling for critic adaptation, posterior-guided exploration, and prior moving average that maintains competitive performance while providing formal certificates.
Several limitations warrant consideration. Our approach requires accurate mixing time estimation (underestimation leads to overconfident bounds), faces computational overhead that may limit scalability to large networks, and uses KL divergence which, despite its analytical convenience, is unstable when distributions diverge significantly and does not respect the geometry of the parameter space (Viallard et al., 2023). Future work could explore efficient posterior approximations, meaningful representations (layer/parameter-level rather than full network posteriors), adaptive mixing time estimation, Wasserstein distance alternatives to KL (Amit et al., 2022; Haddouche & Guedj, 2023; Viallard et al., 2023), and extensions beyond actor-critic frameworks. Despite these constraints, our results believably establish PAC-Bayesian computations for RL as a viable approach for certified performance in modern deep reinforcement learning, particularly valuable in high-stakes applications requiring theoretical guarantees.
@@ B.1 Bounded-differences property for MDP trajectories
As defined in equation 8 and equation 9, our empirical and expected losses are:
@@ B.2 Quantifying the impact of perturbed transitions
Suppose we replace a single transition at position $h$ in trajectory $j$. This perturbation alters the trajectory $\xi^{(j)}$ to $\bar{\xi}^{(j)}$ and effectively changes the cumulative importance weight associated with that trajectory.
By our assumption in Section 3.1, the weighted rewards are effectively bounded in $0,R_{\max}]$. Therefore, we can bound the sensitivity of the empirical loss by treating the weighted signal similarly to the on-policy case. The maximum change in the weighted discounted return due to a perturbation at step $h$ is bounded by the propagation of the error through the discount factor:
@@ B.4 Full accounting of perturbation propagation effects
When divided by $T$ (because $\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)$ averages over $T$ trajectories), we get exactly the same quantity as the sum of coefficients above. Therefore, the bounded-differences condition holds with equality, confirming that our derivation of $\|c\|^{2}$ fully accounts for all propagation effects.
@@ B.5.1 From McDiarmid to moment generating function
McDiarmid’s inequality for Markov chains (inequality 19) provides a concentration inequality on the deviation between empirical and expected losses. From this, we can derive a bound on the moment generating function (MGF) as shown by Paulin (2018):
@@
$\mathbb{E}_{\mathfrak{D}\sim\mathcal{M}^{(T)}}\left[\exp\left(\kappa(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))\right)\right]\leq\exp\left(\frac{\kappa^{2}\|c\|^{2}\tau_{\min}}{8}\right)$ (20)
@@
where $\tau_{\min}$ is the mixing time of the Markov chain induced by a behavior policy $\pi_{b}$.
B.5.2 Final bound via Sub-Gaussian Quadratic Moment
Instead of optimizing the free parameter $\kappa$ in equation 42, which would require a data-independent grid search and a union bound (Alquier, 2024), we leverage the sub-Gaussian nature of our generalization error to derive a tighter bound directly.
From Lemma B.1 (Equation 20), we established that the generalization error $\Delta(\theta)=\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)$ satisfies the sub-Gaussian property with variance proxy $\sigma_{proxy}^{2}$:
$\mathbb{E}_{\mathfrak{D}}\left[\exp(\kappa\Delta(\theta))\right]\leq\exp\left(\frac{\kappa^{2}\|c\|^{2}\tau_{\min}}{8}\right)=\exp\left(\frac{\kappa^{2}\sigma_{proxy}^{2}}{2}\right),\quad\text{where }\sigma_{proxy}^{2}=\frac{\|c\|^{2}\tau_{\min}}{4}.$ (21)
We utilize the property that centered sub-Gaussian random variables satisfy a bound on their quadratic exponential moment. Specifically, following Wainwright (2019) (Theorem 2.6.IV) and the correction by Hellström & Durisi (2020; 2021), for a $\nu$-sub-Gaussian variable $X$, and for any $\lambda\in[0,1)$, we have $\mathbb{E}[\exp(\frac{\lambda X^{2}}{2\sigma_{proxy}^{2}})]\leq\frac{1}{\sqrt{1-\lambda}}$. Choosing $\lambda=\frac{1}{2}$ for simplicity:
$\mathbb{E}_{\mathfrak{D}}\left[\exp\left(\frac{\Delta(\theta)^{2}}{4\sigma_{proxy}^{2}}\right)\right]\leq\sqrt{2}.$ (22)
We now apply the PAC-Bayesian change of measure (Lemma A.2) to the function $f(\theta)=\frac{\Delta(\theta)^{2}}{4\sigma_{proxy}^{2}}$:
@@ B.5.2 Final bound via Sub-Gaussian Quadratic Moment
$\mathbb{E}_{\theta\sim\rho}\left[\frac{(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))^{2}}{4\sigma_{proxy}^{2}}\right]\leq\mathrm{KL}(\rho\|\mu)+\ln\mathbb{E}_{\theta\sim\mu}\left[\exp\left(\frac{\Delta(\theta)^{2}}{4\sigma_{proxy}^{2}}\right)\right].$ (23)
Taking the expectation with respect to the data distribution $\mathfrak{D}\sim\mathcal{M}^{(T)}$ and applying Jensen’s inequality to the concave function $\ln(\cdot)$ followed by Fubini’s theorem to swap expectations in the second term:
$\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}\left[\frac{(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))^{2}}{4\sigma_{proxy}^{2}}\right]$ $\leq\mathbb{E}_{\mathfrak{D}}[\mathrm{KL}(\rho\|\mu)]+\ln\mathbb{E}_{\theta\sim\mu}\mathbb{E}_{\mathfrak{D}}\left[\exp\left(\frac{\Delta(\theta)^{2}}{4\sigma_{proxy}^{2}}\right)\right]$ (24)
[…]
Multiplying by $4\sigma_{proxy}^{2}$ and substituting $\sigma_{proxy}^{2}=\frac{\|c\|^{2}\tau_{\min}}{4}$:
$\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}\left[(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))^{2}\right]$ $\leq\|c\|^{2}\tau_{\min}\left(\mathbb{E}_{\mathfrak{D}}[\mathrm{KL}(\rho\|\mu)]+\ln\sqrt{2}\right).$ (26)
Finally, we apply Jensen’s inequality to the left-hand side, using $(\mathbb{E}[X])^{2}\leq\mathbb{E}[X^{2}]$:
$\left(\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}\left[\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)\right]\right)^{2}\leq\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}\left[(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))^{2}\right].$ (27)
Taking the square root gives the final in-expectation bound:
$\left|\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}[\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)]\right|\leq\sqrt{\|c\|^{2}\tau_{\min}\left(\mathbb{E}_{\mathfrak{D}}[\mathrm{KL}(\rho\|\mu)]+\frac{1}{2}\ln 2\right)}.$ (28)
B.5.3 High-probability bound
To obtain the high-probability result (Theorem 3.2), we apply Markov’s inequality directly to the random variable $Z=\exp\left(\frac{\Delta(\theta)^{2}}{4\sigma_{proxy}^{2}}\right)$ within the change of measure step, similar to Hellström & Durisi (2021, Eq. 3). With probability at least $1-\delta$:
$\mathbb{E}_{\theta\sim\rho}\left[(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))^{2}\right]\leq 4\sigma_{proxy}^{2}\left(\mathrm{KL}(\rho\|\mu)+\ln\frac{\sqrt{2}}{\delta}\right).$ (29)
Using Jensen’s inequality $(\mathbb{E}_{\theta\sim\rho}[\Delta(\theta)])^{2}\leq\mathbb{E}_{\theta\sim\rho}[\Delta(\theta)^{2}]$ and taking the square root:
$\left|\mathbb{E}_{\theta\sim\rho}[\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)]\right|\leq\sqrt{\|c\|^{2}\tau_{\min}\left(\mathrm{KL}(\rho\|\mu)+\ln\frac{\sqrt{2}}{\delta}\right)}.$ (30)
Substituting $\|c\|^{2}$ from equation 12 yields the result in Theorem 3.2.
$\mathbb{E}_{\theta\sim\rho}[V_{\pi_{\theta}}]\geq\mathbb{E}_{\theta\sim\rho}[-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)]-\sqrt{\|c\|^{2}\tau_{\min}\left(\mathrm{KL}(\rho\|\mu)+\ln\frac{\sqrt{2}}{\delta}\right)}.$ (31)
where he true expected value is lower-bounded by an empirical estimate minus an uncertainty term that accounts for limited data ($1/T$), temporal correlations ($\tau_{\min}$), and posterior complexity ($\mathrm{KL}(\rho\|\mu)$).
B.5.4 Optimization of the bound
The high-probability bound derived above takes the following form:
$\mathbb{E}_{\theta\sim\rho}[\mathcal{L}(\theta)]\leq\underbrace{\mathbb{E}_{\theta\sim\rho}[\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)]}_{\text{Linear in }\rho}+\underbrace{\sqrt{\|c\|^{2}\tau_{\min}\left(\mathrm{KL}(\rho\|\mu)+\ln\frac{\sqrt{2}}{\delta}\right)}}_{\text{Concave in KL (Non-convex in }\rho)}.$ (32)
Directly optimizing this objective with respect to the posterior $\rho$ is challenging because the square root of the KL divergence makes the overall objective non-convex. To address this, we employ a variational approximation to obtain a surrogate objective that is amenable to alternating convex optimization.
We utilize the algebraic identity for the square root function: for any $x\geq 0$, $\sqrt{x}=\inf_{\lambda>0}\left(\frac{x}{2\lambda}+\frac{\lambda}{2}\right)$. By identifying $x$ with the complexity term inside the square root, we obtain an upper bound on the true risk that depends on the auxiliary parameter $\lambda$:
$\mathbb{E}_{\theta\sim\rho}[\mathcal{L}(\theta)]\leq\inf_{\lambda>0}\mathcal{J}(\rho,\lambda),$ (33)
where the surrogate objective $\mathcal{J}(\rho,\lambda)$ is defined as:
$\mathcal{J}(\rho,\lambda)=\mathbb{E}_{\theta\sim\rho}[\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)]+\frac{\|c\|^{2}\tau_{\min}\left(\mathrm{KL}(\rho\|\mu)+\ln\frac{\sqrt{2}}{\delta}\right)}{2\lambda}+\frac{\lambda}{2}.$ (34)
This surrogate objective has two desirable properties that justify the alternating optimization procedure in our algorithm (Section 4):
Convexity in $\rho$: For a fixed $\lambda$, the term becomes a linear scaling of the KL divergence, which is convex in $\rho$. This allows for stable gradient-based updates of the posterior parameters.
Convexity in $\lambda$: For a fixed $\rho$, the function is strictly convex in $\lambda$, allowing for an analytic solution:
$\lambda^{*}=\sqrt{\|c\|^{2}\tau_{\min}\left(\mathrm{KL}(\rho\|\mu)+\ln\frac{\sqrt{2}}{\delta}\right)}.$ (35)
Minimizing $\mathcal{J}(\rho,\lambda)$ via alternating updates is therefore mathematically equivalent to minimizing the rigorous high-probability bound derived in Section B.5.3.
Appendix C Alternative derivation via union bound and a grid over $\kappa$
Below, we provide another alternative derivation, keeping the parameter $\kappa$ and optimizing it via a grid search and union bound following the procedure from Alquier (2024).
C.0.1 PAC-Bayes change of measure
Starting from the MGF bound above (Lemma B.1), we can follow the standard PAC-Bayes derivation. Let $\Theta$ be our parameter space and let $\mu\in\mathcal{P}(\Theta)$ be a prior distribution over $\Theta$ chosen independently of the data. For any posterior distribution $\rho\in\mathcal{P}(\Theta)$ (which may depend on $\mathfrak{D}$), we apply the change of measure inequality that follows from Donsker–Varadhan variational formula (Donsker & Varadhan, 1975; 1983).
Let $f(\theta)=\kappa(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))$. Applying Lemma A.2:
$\mathbb{E}_{\theta\sim\rho}[\kappa(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))]$ $\leq\mathrm{KL}(\rho\|\mu)+\ln\mathbb{E}_{\theta\sim\mu}[\exp(\kappa(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)))]$ (36)
C.0.2 Combining with the MGF bound
@@
$\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}[\kappa(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))]$ $\leq\mathbb{E}_{\mathfrak{D}}\left[\mathrm{KL}(\rho\|\mu)\right]+\mathbb{E}_{\mathfrak{D}}\left[\ln\mathbb{E}_{\theta\sim\mu}[\exp(\kappa(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)))]\right]$ (37)
Applying Jensen’s inequality to the concave function $\ln(\cdot)$ yields:
$\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}[\kappa(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))]$ $\leq\mathbb{E}_{\mathfrak{D}}\left[\mathrm{KL}(\rho\|\mu)\right]+\ln\mathbb{E}_{\mathfrak{D}}\left[\mathbb{E}_{\theta\sim\mu}[\exp(\kappa(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)))]\right]$ (38)
By Fubini-Tonelli’s theorem, we swap the order of expectations in the second term and apply the MGF bound from Lemma B.1:
$\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}[\kappa(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta))]$ $\leq\mathbb{E}_{\mathfrak{D}}\left[\mathrm{KL}(\rho\|\mu)\right]+\ln\mathbb{E}_{\theta\sim\mu}\mathbb{E}_{\mathfrak{D}}[\exp(\kappa(\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)))]$ (39)
$\leq\mathbb{E}_{\mathfrak{D}}\left[\mathrm{KL}(\rho\|\mu)\right]+\ln\mathbb{E}_{\theta\sim\mu}\left[\exp\left(\frac{\kappa^{2}\|c\|^{2}\tau_{\min}}{8}\right)\right]$ (40)
$=\mathbb{E}_{\mathfrak{D}}\left[\mathrm{KL}(\rho\|\mu)\right]+\frac{\kappa^{2}\|c\|^{2}\tau_{\min}}{8}$ (41)
@@
$\mathbb{E}_{\mathfrak{D}}\mathbb{E}_{\theta\sim\rho}[\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)]$ $\leq\frac{\mathbb{E}_{\mathfrak{D}}\left[\mathrm{KL}(\rho\|\mu)\right]}{\kappa}+\frac{\kappa\|c\|^{2}\tau_{\min}}{8}$ (42)
@@
Appendix D A note on the Markov assumption for Bellman errors